# A Multimodal Data Fusion Generative Adversarial Network for Real Time Underwater Sound Speed Field Construction

### v1

@@ 3.3.2 Accuracy in Deep and Shallow Water
Figure 5: Comparison of sound speed profiles at different locations and depths. Each row represents a specific depth (200 m, 400 m, 1000 m, and 1975 m), while each column represents a specific location (Location 1: ($49.5^{\circ}$, $13.5^{\circ}$); Location 2: ($49.5^{\circ}$, $28.5^{\circ}$); Location 3: ($52.5^{\circ}$, $25.5^{\circ}$)). The figures illustrate the comparison of sound speed profile predictions using four different methods: MDF-RAGAN, CNN, SITP, and MEAN at each location and depth.
@@
Table 3: Comparison of RMSE in shallow sea areas at different depths
@@
MDF-RAGAN ($49.5^{\circ}$, $16.5^{\circ}$) 0.2481 0.2191 0.1812
($49.5^{\circ}$, $34.5^{\circ}$) 0.2561 0.2376 0.1280
($49.5^{\circ}$, $28.5^{\circ}$) 0.2641 0.2340 0.2004
($49.5^{\circ}$, $31.5^{\circ}$) 0.1856 0.1666 0.1280
CNN ($49.5^{\circ}$, $16.5^{\circ}$) 0.4843 0.4618 0.3971
($49.5^{\circ}$, $34.5^{\circ}$) 0.4346 0.4835 0.2941
($49.5^{\circ}$, $28.5^{\circ}$) 0.4994 0.4754 0.3720
($49.5^{\circ}$, $31.5^{\circ}$) 0.4134 0.3883 0.2941
MEAN ($49.5^{\circ}$, $16.5^{\circ}$) 0.3000 0.3239 0.2662
($49.5^{\circ}$, $34.5^{\circ}$) 1.4430 1.5012 0.5114
($49.5^{\circ}$, $28.5^{\circ}$) 0.7325 0.5776 0.4130
($49.5^{\circ}$, $31.5^{\circ}$) 0.9546 0.7745 0.5114
SITP ($49.5^{\circ}$, $16.5^{\circ}$) 0.2550 0.2363 0.1848
($49.5^{\circ}$, $34.5^{\circ}$) 1.0823 1.1221 0.4604
($49.5^{\circ}$, $28.5^{\circ}$) 0.5201 0.4192 0.2921
($49.5^{\circ}$, $31.5^{\circ}$) 0.8164 0.6978 0.4604
The shallow sea profile enters the scattering layer at the stage of 200-1000 m. The sound velocity changes in this region are smoother, but there are still small oscillations. Although the errors of MEAN and SITP are slightly reduced, MDF-RAGAN is still much better than the other methods, and the prediction curves of CNN are alleviated at this layer, but it is still difficult to maintain the smoothness at the full depth. In contrast, MDF-RAGAN successfully smoothes out the high-frequency noise and closely follows the measured curves by virtue of the cross-modal attention that dynamically adjusts the fusion weights of the neighborhood and labeled features at each layer of depth. More detailed results at 4 typical locations are given in Table 3.
@@
3.4 Discussions
@@
The comprehensive comparison results based on the whole-profile mean sound speed error (RMSE) are summarized in Table 4. It can be seen that MDF-RAGAN not only outperforms CNN and SITP by nearly a factor of two, but also achieves about 65.8% RMSE reduction compared to MEAN, which fully reflects the enhancement of overall profile matching by multi-source fusion and cross-modal attention.
Table 4: Comparison of overall RMSE (m/s) across different methods
Methods RMSE (m/s) Parameters Inference Time (ms)
MDF-RAGAN 0.1479 7.7M 4.48
CNN 0.3107 4.69M 0.87
SITP 0.3384 - -
MEAN 0.4335 - -

### v3

@@ III-C3 Accuracy in Deep and Shallow Water
Figure 4: Comparison of sound speed profiles at different locations and depths. Each row represents a specific depth (200 m, 400 m, 1000 m, and 1975 m), while each column represents a specific location (Location 1: ($49.5^{\circ}$S, $13.5^{\circ}$E); Location 2: ($49.5^{\circ}$S, $28.5^{\circ}$E); Location 3: ($52.5^{\circ}$S, $25.5^{\circ}$E)). The figures illustrate the comparison of sound speed profile predictions using four different methods: MDF-RAGAN, CNN, SITP, and MEAN at each location and depth.
@@
The shallow sea profile enters the scattering layer at the stage of 200-1000 m. The sound velocity changes in this region are smoother, but there are still small oscillations. Although the errors of MEAN and SITP are slightly reduced, MDF-RAGAN is still much better than the other methods, and the prediction curves of CNN are alleviated at this layer, but it is still difficult to maintain the smoothness at the full depth. In contrast, MDF-RAGAN successfully smoothes out the high-frequency noise and closely follows the measured curves by virtue of the cross-modal attention that dynamically adjusts the fusion weights of the neighborhood and labeled features at each layer of depth. More detailed results at 4 typical locations are given in Table V.
TABLE V: Comparison of RMSE in shallow sea areas at different depths
@@
MDF-RAGAN ($49.5^{\circ}$S, $16.5^{\circ}$E) 0.248 0.219 0.181
($49.5^{\circ}$S, $34.5^{\circ}$E) 0.256 0.237 0.128
($49.5^{\circ}$S, $28.5^{\circ}$E) 0.264 0.234 0.200
Att-CNN ($49.5^{\circ}$S, $16.5^{\circ}$E) 0.484 0.462 0.397
($49.5^{\circ}$S, $34.5^{\circ}$E) 0.435 0.483 0.294
($49.5^{\circ}$S, $28.5^{\circ}$E) 0.499 0.475 0.372
MEAN ($49.5^{\circ}$S, $16.5^{\circ}$E) 0.300 0.324 0.266
($49.5^{\circ}$S, $34.5^{\circ}$E) 1.443 1.501 0.511
($49.5^{\circ}$S, $28.5^{\circ}$E) 0.733 0.578 0.413
SITP ($49.5^{\circ}$S, $16.5^{\circ}$E) 0.255 0.236 0.185
($49.5^{\circ}$S, $34.5^{\circ}$E) 1.082 1.122 0.460
($49.5^{\circ}$S, $28.5^{\circ}$E) 0.520 0.419 0.292
@@
III-C4 Ablation Study
To further quantify the contribution of each core component in MDF-RAGAN, we conduct an ablation study by progressively disabling different modules while keeping the remaining settings unchanged. Specifically, we consider four degraded variants: (i) w/o residual, which removes the residual perturbation branch and directly regresses the full SSP instead of the perturbation around the neighbourhood mean profile; (ii) w/o FMB, which discards the Feature Mapping Block and thus removes the depth-wise enhancement of SSP features; (iii) w/o SST, which feeds only historical SSPs and geographic coordinates into the network while entirely removing the SST modality, thereby breaking the explicit coupling between surface temperature and subsurface sound speed; and (iv) w/o CMPAB, which eliminates the Cross-Modal Perturbation Attention Block so that the model can no longer adaptively fuse labels and reference SSPs across large-scale spatial regions.
TABLE VI: Ablation study on the main components of MDF-RAGAN.
Variant MDF-RAGAN w/o residual w/o FMB w/o SST w/o CMPAB
RMSE (m/s) 0.148 1.006 0.256 0.641 0.675
As reported in Table VI, removing the residual perturbation formulation (w/o residual) severely degrades the reconstruction accuracy, with RMSE increasing from 0.148 m/s to 1.006 m/s, which confirms that learning fine-grained perturbations around the neighborhood mean profile is much easier than regressing the entire SSP from scratch. The performance drops observed in the w/o FMB and w/o CMPAB variants further demonstrate that both depth-wise feature enhancement and cross-modal perturbation attention are indispensable for capturing thermocline structures and long-range spatial dependencies. In particular, the w/o SST variant yields a RMSE of 0.641 m/s, highlighting that explicitly incorporating SST as an additional modality provides crucial physical information on upper-ocean thermal gradients, and plays a key role in improving the fidelity of SSP prediction. From a physical perspective, SST controls the upper-ocean thermohaline gradient and thus provides a strong prior for the vertical structure of the SSP, especially around the thermocline. By explicitly incorporating SST as a remote-sensing modality, MDF-RAGAN can exploit globally available satellite products to improve SSP estimation in data-sparse regions where in-situ profiles are scarce.
III-D Discussions
@@
TABLE VII: Comparison of overall RMSE (m/s), model size, inference time and training time across different methods
Methods RMSE (m/s) Parameters Inference Time (ms) Training Time
MDF-RAGAN 0.148 7.7M 4.48 1 h 18 min 2 s
Att-CNN 0.311 4.69M 0.87 38 min 49 s
SITP 0.338 - - -
MEAN 0.434 - - -
@@
The comprehensive comparison results based on the whole-profile mean sound speed error (RMSE) are summarized in Table VII. It can be seen that MDF-RAGAN not only outperforms CNN and SITP by nearly a factor of two, but also achieves about 65.8% RMSE reduction compared to MEAN, which fully reflects the enhancement of overall profile matching by multi-source fusion and cross-modal attention. By comparing the scale of model parameters and the execution time, our proposed model occupies more storage and requires more computation time. However, considering that training can be completed offline in advance and the execution time is still in milliseconds, it is acceptable in practical applications.
@@ IV Conclusion
Acknowledgment
SSP data were taken from the GDCSM_Argo gridded Argo dataset [26], and the SST data were obtained from the NOAA OISST monthly average product [27]. Data will be available on request. The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.